CAN SOMEONE PLEASE HELP ME AND SHOW ME HOW TO WORK OUT THESE PROBLEMS.I IF U BREAK IT DOWN FOR ME I WILL UNDERSTAND IT BETTER.
SKETCH A GRAPH OF Y=(X/2)-1. INCLUDE THE SLOPE AND BOTH INTERCEPTS.

THANK YOU!

Answer by mananth(16946)

Y=(X/2)-1


Y intercept plug x=0
when x= 0 y= -1 ( 0 , -1 )
X intercept plug y=0
when y = 0 x= 2 ( 2 , 0 )